OpenStudy (anonymous):

@wendi20 Lets make another page, because the other one is long 0.7z+1.9+0.1z=5.5-0.4z

OpenStudy (anonymous):

@wendi20 can you show me your own problem solving again

OpenStudy (anonymous):

answer was 30

OpenStudy (anonymous):

you meant z=3 not z=30

OpenStudy (anonymous):

0.7z+1.9+0.1z=5.5-0.4z 0.8z+1.9=5.5-0.4z 0.8z+0.4z+1.9=5.5 .12z=5.5-1.9 .12z=3.6 then div 3.6/.12

OpenStudy (anonymous):

we didnt move decimals right

OpenStudy (anonymous):

yes so its \[\frac{ 3.6 }{ 1.2 }\]

OpenStudy (anonymous):

are you using calculator? if you have one, try it?

OpenStudy (anonymous):

1/8(5y+64)=1/4(20+2y) hld sec let me see if I can do it if you dont mind

OpenStudy (anonymous):

do you know whats the first thing you can do here right?

OpenStudy (anonymous):

parathesis first

OpenStudy (anonymous):

once taken out to there signs change like neg and pos?

OpenStudy (anonymous):

DISTRIBUTIVE FIRST

OpenStudy (anonymous):

meaning sending the 5y over to the other side and it would then be a neg 5y

OpenStudy (anonymous):

Nope xD meaning sending the 1/8 over to the parenthesis and do the same thing to the other side

OpenStudy (anonymous):

-1/4(5y+64)=-1/8(20+2y)

OpenStudy (anonymous):

or (5y+64)=-1/8+1/4(20+2y)

OpenStudy (anonymous):

No again.. \[\frac{ 1 }{ 8 }(5y+64)=\frac{ 1 }{ 4 }(20+2y)\]  can you solve this first